If I Had walked 1 Km/h faster,I would have taken 10 minutes less to walk 2 km.Find the speed of my walking?
Answer by amoresroy(361)

If I Had walked 1 Km/h faster,I would have taken 10 minutes less to walk 2 km.Find the speed of my walking?

Let s = speed of walking
2/s = travel time at speed s
2/(s+1)= travel time at speed (s+1)
2/s - 2/(s+1) = 1/6
Simply equation
s^2 + s -12 = 0
(s+4)(s-3) = 0
s-3 = 0
s=3
The speed of walking is 3 km/h